# CONQURE: A Co-Execution Environment for Quantum and Classical Resources

### v2

@@ Abstract
Index Terms:
Quantum Computing, Quantum Cloud Portal, Job Management, Domain-Specific Languages
@@ I Introduction
Quantum computing has the potential to disrupt classical computing for a number of fields. These fields include clinical research [1], optimization [2], high-energy physics [3], finance [4] and logistics [5]. The catalog of problems where quantum computing promises potential for significant speedup over its classical counterparts is only increasing. As such, quantum computing has reached an early (yet still experimental) state of maturity. A number of device technologies ranging from superconducting devices (e.g., IBM Q, Rigetti, QCI, OQC, Google, Amazon) to ion traps (IonQ, Quantinuum) to neutral atoms (QuEra, Pasqal) have become available through cloud access, e.g., via Amazon Braket [6], Azure Quantum [7], qBraid [8] or directly from the vendors. Existing commercial solutions feature quantum computing ecosystems on the basis of python packages to facilitate software development over a number of hardware devices. Today’s leading quantum computing hardware technologies include Superconducting qubits (IBM, Google, Amazon, Alice & Bob), Ion-Traps (IONQ, Quantinuum) and Neutral Atoms (QuEra Computing Inc.). Software libraries like Qiskit [9], Cirq [10], Tket [11] and Pennylane [12] enable users to design quantum workloads while lower-level control of the hardware is enabled by interfacing through proprietary software layers, or research infrastructure such as DAX [13], ARTIQ [14], Qubic [15, 16], and QICK [17]. This ecosystem of hardware and software technologies supports a growing interest in quantum computing and is evolving rapidly.
@@ III-B2 Scalability and Performance
Given that CONQURE is intended for use by researchers across labs, scalability was a key consideration in the API design. The system must be able to handle high volumes of requests without compromising performance. Fig 2 outlines the flow of data across different parts of the architecture. Our API supports asynchronous job submissions, allowing users to submit multiple jobs without needing to wait for their completion. This is particularly important here because job execution times are expected to vary significantly. SLURM ensures efficient resource utilization with job queuing that can be tailored for each hardware device.
@@ III-B4 Modularity and Adaptability
Listing LABEL:lst:qpu demonstrates repeated execution using our proposed quantum offloading strategy. The OpenMP-Q framework enables reverse offloading, where the quantum kernel remains active while classical computation runs asynchronously on the host. This reduces task creation overhead and allows classical computations (e.g., parameter updates) to directly influence subsequent quantum gate operations.
More specifically, results from one iteration of the quantum kernel may trigger classical code execution (e.g., a solver), whose output is relayed back to the quantum processor. These values are then incorporated into the subsequent quantum gate executions, such as phase-angle adjustments. This approach can be realized through on-the-fly parametric pulse shaping within the FPGA that controls the quantum device.
To facilitate quantum-host communication, we utilize a quantum class object (see Sec. IV) to copy qubits into classical space, enabling their communication with other MPI nodes during parallel execution. This ensures that classical solving itself becomes parallelized, while the quantum kernel continues execution with updated angle values.
Furthermore, the classical solver could be implemented as a GPU kernel or distributed across CPU cores on each node. The computed results are aggregated to determine the best angle parameters before the next quantum kernel iteration. MPI parallelization is optional; if MPI calls are omitted, execution remains limited to a single node. However, with MPI support, our model extends to multi-QPU execution enabling result consolidation across quantum processors.
@@ V-A CONQURE Cloud Queue
% Total execution time when considering \$n\$ circuits executed on the same quantum device with multi-execution and without. (graph)
% $\begin{itemize}$
[…]
% Fidelity threshold vs Reduction in Execution Time (graph)
% $\begin{itemize}$
[…]
% Probability of Successful Trial, PST (graph)
@@ Acknowledgment
This work was supported in part by NSF awards MPS-2410675, PHY-1818914, PHY-2325080, MPS-2120757, and CISE-2316201. We would like to thank Andrew Van Horn from Duke Quantum Center for his support in the testing of CONQURE on real devices.

### v3

@@ I Introduction
Quantum computing has the potential to disrupt classical computing for a number of fields. These fields include clinical research [1], optimization [2], high-energy physics [3], finance [4] and logistics [5]. The catalog of problems where quantum computing promises potential for significant speedup over its classical counterparts is only increasing. As such, quantum computing has reached an early (yet still experimental) state of maturity. A number of device technologies ranging from superconducting devices (e.g., IBM Q, Rigetti, QCI, OQC, Google, Amazon) to ion traps (IonQ, Quantinuum) to neutral atoms (QuEra, Pasqal) have become available through cloud access, e.g., via Amazon Braket [6], Azure Quantum [7], qBraid [8] or directly from the vendors. Existing commercial solutions feature quantum computing ecosystems on the basis of python packages to facilitate software development over a number of hardware devices. Today’s leading quantum computing hardware technologies include Superconducting qubits (IBM, Google, Amazon, Alice & Bob), Ion-Traps (IONQ, Quantinuum) and Neutral Atoms (QuEra Computing Inc.). Software libraries like Qiskit [9], Cirq [10], Tket [11], Pennylane [12] and CUDA-Q [13] enable users to design quantum workloads while lower-level control of the hardware is enabled by interfacing through proprietary software layers, or research infrastructure such as DAX [14], ARTIQ [15], Qubic [16, 17], and QICK [18]. This ecosystem of hardware and software technologies supports a growing interest in quantum computing and is evolving rapidly.
@@ III-B2 Scalability and Performance
Given that CONQURE is intended for use by researchers across labs, scalability was a key consideration in the API design. The system must be able to handle high volumes of requests without compromising performance. Fig 2 outlines the flow of data across different parts of the architecture. Our API supports asynchronous job submissions, allowing users to submit multiple jobs without needing to wait for their completion. This is particularly important here because job execution times are expected to vary significantly. SLURM [30] ensures efficient resource utilization with job queuing that can be tailored for each hardware device.
@@ III-B4 Modularity and Adaptability
[…]
proposed quantum offloading strategy. The OpenMP-Q framework enables
reverse offloading, where the quantum kernel remains active while
classical computation runs asynchronously on the host. This reduces
task creation overhead and allows classical computations (e.g.,
parameter updates) to directly influence subsequent quantum gate
operations.
More specifically, results from one iteration of the quantum kernel
may trigger classical code execution (e.g., a solver), whose output is
relayed back to the quantum processor. These values are then
incorporated into the subsequent quantum gate executions, such as
phase-angle adjustments. This approach can be realized through
on-the-fly parametric pulse shaping within the FPGA that controls the
quantum device.
To facilitate quantum-host communication, we utilize a quantum class
[…]
enabling their communication with other MPI nodes during parallel
[…]
parallelized, while the quantum kernel continues execution with
updated angle values.
Furthermore, the classical solver could be implemented as a GPU kernel
or distributed across CPU cores on each node. The computed results are
aggregated to determine the best angle parameters before the next
quantum kernel iteration. MPI parallelization is optional; if MPI
calls are omitted, execution remains limited to a single
node. However, with MPI support, our model extends to multi-QPU
execution enabling result consolidation across quantum processors.
@@ Acknowledgment
This work was supported in part by NSF awards MPS-2410675, PHY-1818914, PHY-2325080, OMA-2120757, and CISE-2316201. We would like to thank Andrew Van Horn from Duke Quantum Center for his support in the testing of CONQURE on real devices.